OpenStudy (anonymous):

the equation of the normal line at x=1

OpenStudy (jonnyvonny):

f(x)=x^2-3 f*(x)=2x

OpenStudy (anonymous):

y=mx+b

OpenStudy (anonymous):

-2=-1/2(1)+b

OpenStudy (jonnyvonny):